OpenStudy (anonymous):

DANGER! need help! please? please help! crashing for finals! find cos(u+v) given that sin u=3/5 and cos v=-5/13 (both u and v are in Quad 2

OpenStudy (anonymous):

Use the cosine addition formula:\[\cos(u+v) = \cos u \cos v - \sin u \sin v\]Also, don't forget the Pythagorean identity (so once you know sine you have cosine).\[\sin^2 \theta + \cos^2 \theta = 1\]

OpenStudy (anonymous):

thanks i got this problem already but thank u ! :)